MidnightMoonLight:

Jared bought 5 1/6 pounds of sand and 2 3/4 pounds of gravel. how many more pounds of sand did Jared buy then gravel?

MrMudd183:

what do you think the first thing to do is?

MidnightMoonLight:

@mrmudd183 wrote:
what do you think the first thing to do is?
possibly to subtract the gravel from the sand?

MrMudd183:

Close, what I would do is get the fractions into a common denominator

RaeRae:

Then subtract.

MrMudd183:

Both fractions can have 12 as a common denominator, so what would you do to get both denominators to be 12?

Imagine:

Also, turn the fractions into improper fractions.

Imagine:

Improper Fraction: \[11/6,11/4\] Find the common denominator of 12. \[11/6(2)=22/12\] \[11/4(3)=33/12\]